OpenStudy (anonymous):

Write each expression as the product of two binomials. x2 + 7x + x + 7

OpenStudy (anonymous):

x2 + 7x + x + 7 = x(x+7)+(x+7) = (x+7)(x+1)

jhonyy9 (jhonyy9):

(x+1)(x+7)

OpenStudy (anonymous):

Wait how do you go from x(x+7)+(x+7) to = (x+7)(x+1) does it just mean what multiplied equals 7 ?